# Comparative Evaluation of 3D Reconstruction Methods for Object Pose Estimation

### v1

@@ Abstract
Object pose estimation is essential to many industrial applications involving robotic manipulation, navigation, and augmented reality. Current generalizable object pose estimators, i.e., approaches that do not need to be trained per object, rely on accurate 3D models. Predominantly, CAD models are used, which can be hard to obtain in practice. At the same time, it is often possible to acquire images of an object. Naturally, this leads to the question whether 3D models reconstructed from images are sufficient to facilitate accurate object pose estimation. We aim to answer this question by proposing a novel benchmark for measuring the impact of 3D reconstruction quality on pose estimation accuracy. Our benchmark provides calibrated images for object reconstruction registered with the test images of the YCB-V dataset for pose evaluation under the BOP benchmark format. Detailed experiments with multiple state-of-the-art 3D reconstruction and object pose estimation approaches show that the geometry produced by modern reconstruction methods is often sufficient for accurate pose estimation. Our experiments lead to interesting observations: (1) Standard metrics for measuring 3D reconstruction quality are not necessarily indicative of pose estimation accuracy, which shows the need for dedicated benchmarks such as ours. (2) Classical, non-learning-based approaches can perform on par with modern learning-based reconstruction techniques and can even offer a better reconstruction time-pose accuracy tradeoff. (3) There is still a sizable gap between performance with reconstructed and with CAD models. To foster research on closing this gap, our benchmark is publicly available at github.com/VarunBurde/reconstruction_pose_benchmark.
@@ 1 Introduction
Object pose estimators predict the position and orientation of an object in a given image. Object pose estimation is a prerequisite to various applications including robotic manipulation and virtual tasks such as highlighting individual components of large machinery in AR applications. Existing methods can be categorized into CAD-based methods, e.g., approaches using feature matching [22, 67, 84], templates [41, 40], regression [46, 107, 35], or render-and-compare strategies [58, 56, 55, 93], and CAD-free methods, e.g., methods using pose regression [111, 62, 37], analysis-by-synthesis [19], object coordinate regression [60, 79, 14, 7, 16], feature matching [99], or render-and-compare [78]. CAD-based methods usually lead the BOP benchmark for object pose estimation [45, 47, 2].
@@
The 3D models involved are often generated by CAD applications in complex setups or reconstructed via range scanners [26] that produce highly accurate meshes. Naturally, this process can be time-consuming. Also, in many practical settings, e.g., service robots, it is infeasible to obtain such a 3D model for each object class and object instance that might be encountered. A more practical approach in such scenarios would be to reconstruct 3D object models on the fly from images.
@@
Generalization is another factor limiting the deployment of pose estimators: in the example where a robot faces a new object and generates an on-the-fly 3D model, most of the estimators still require training or finetuning on the new object. This induces a computational overhead not only because of the network training but also because the robot must capture enough views for the training to converge. Instead, deploying CAD-based estimators that generalize to unseen objects [55, 12] is more efficient since it can be used without finetuning and 3D reconstruction can work even with sparse views. Two such estimators have recently set the State-of-the-Art (SotA) with a CAD-based render-and-compare strategy [55, 12].
@@
In this paper, we investigate using 3D reconstructions from RGB images for object pose estimation. In particular, we want to quantify the object pose accuracy gap between using high-quality CAD models and 3D models obtained from images inside generalizable object pose estimators [55, 12]. For this purpose, we propose a new benchmark. Compared to existing benchmarks for 3D reconstruction [90, 53, 48, 102], our benchmark does not treat 3D reconstruction as a task unto itself but rather evaluates the resulting 3D models inside a higher-level task (namely object pose estimation). Thus, we mainly measure 3D reconstruction performance by the accuracy of the estimated poses rather than the accuracy of the resulting 3D models itself. Our benchmark is based on the YCB-V [111] dataset for object pose evaluation: it is made of 21 objects selected from the YCB dataset [17], including small objects, objects with low texture, complex shapes, high reflectance, and multiple symmetries. For each object, our benchmark consists of two components: (1) Images of the object captured by a robot arm that can be used for 3D reconstruction. We provide image sets of different sizes as, in certain applications, data capture and 3D reconstruction times can be important. Providing smaller subsets of images allows us to simulate such scenarios. (2) A set of test images for object pose estimation together with ground truth poses registered with the image sets from (1). Compared to the data for the first component, which we captured ourselves, we use the test images provided by the YCB-V [111] dataset for the second component. Given the two components, we evaluate multiple state-of-the-art 3D reconstruction methods, including both classical methods based on multi-view stereo [88, 82] and recent learning-based approaches using neural implicit representations [113, 116, 114, 75, 109]. Our experiments show that standard metrics for 3D reconstruction quality (both geometry-based and appearance-based) do not necessarily directly correlate with object pose estimation accuracy. This clearly shows a need for benchmarks such as ours that directly measure the impact of 3D reconstruction algorithms on object pose estimation’s accuracy.
@@
In summary, this paper makes the following contributions: (1) We propose a new benchmark to evaluate 3D reconstruction algorithms in the context of object pose estimation. Our benchmark aims to measure to what degree 3D models reconstructed from images instead of CAD models can be used for object pose estimation. (2) We evaluate a wide range of state-of-the-art 3D reconstruction algorithms on our benchmark in terms of object pose accuracy, data requirements to reach high accuracy, and reconstruction times. The most important insights are: (a) There is a clear need for benchmarks that evaluate 3D reconstruction algorithms in the context of object pose estimation. (b) Classical, non-learning-based approaches can perform similar to modern learning-based reconstruction techniques while offering faster reconstruction times. (c) The top-performing reconstruction methods are the same for both object pose estimators considered in our benchmark [55, 12], suggesting that all object pose estimators might potentially benefit from progress in the area of 3D reconstruction. (d) While it is possible to obtain a similar performance as with CAD models for some types of objects, there is still a sizable gap in performance for other types of objects, especially for objects with fine details and reflective surfaces. Clearly, the problem of 3D object reconstruction is not yet solved. (3) The benchmark data is available at github.com/VarunBurde/reconstruction_pose_benchmark.
@@ 2 Related Work
Object pose estimation.  The main approaches include feature matching [27, 42, 22, 67, 84, 66], template matching [39, 110], pose regression [111, 62, 37], 3D-coordinate-regression [46, 107, 35, 13, 14, 60, 79, 103, 16], and render-and-compare [58, 56, 55, 93]. One key property for the deployment of pose estimators is their generalization that falls into three categories: no generalization, generalization to object categories seen at training time [61, 108, 19, 66], and generalization to any object [55, 12, 6, 73, 76, 70]. Pose estimators that generalize to any objects are the most suitable in practical scenarios as it is unrealistic to train the estimators on any object instance or category that might be encountered during deployment. However, they often rely on CAD models and it is also impractical to generate CAD models for all objects beforehand. This prompts recent contributions on generalizable pose estimators to use 3D models generated on the fly for the pose estimation [18, 99, 36, 78, 73, 12]. At the time of the writing, FoundationPose [12] sets the state-of-the-art which motivates its integration in the benchmark to evaluate the 3D reconstructions. Whenever a CAD model is not available, it generates a 3D model and use it in a render-and-compare fashion. We also evaluate the 3D reconstructions with Megapose [55], another SotA render-and-compare method that operates on RGB images.
@@
3D Reconstruction.  Given a set of calibrated images with known camera poses, Multi-View Stereo (MVS) generates a dense 3D reconstruction of the scene with explicit representations such as dense point clouds [31, 57], depth maps [97, 88, 11, 50] surface hulls [32, 94], or voxels [91, 54, 95, 105]. One advantage of depth-based approaches is the recovery of fine details at a reasonable memory budget contrary to surface or voxel based methods which memory requirements usually increase with the resolution of the 3D reconstruction. This motivates the benchmark to focus on depth-based methods with the open-source COLMAP [87, 88] and its industrial counterpart Reality Capture [82].
@@
Alternatives to explicit reconstructions are implicit representations [29, 23, 72], i.e., continuous functions defined over the 3D space that embed properties about the scene. An example is the Signed Distance Function (SDF) that outputs the distance between a given 3D point and the closest surface. Defining implicit functions is often not tractable so they are instead approximated, often with neural networks [68, 77, 109, 96]. With the advent of differentiable rendering, several implicit functions relevant for 3D reconstruction can be learned simply from calibrated-image-supervision including volumetric radiance field [69], occupancy network [75], and SDF [112, 74]. Volumetric methods [69, 63] represent the scene as a colored volume where a network outputs a volume density and a color for each point in the space. Rendering the volume density produces depth maps suitable to generate 3D meshes using Poisson surface resconstruction [52]. Volumetric methods are optimized for novel view synthesis only whereas optimizing SDF networks involves more geometric constraints [34] to ensure that the network exhibits SDF properties such as 0 values on the surface and smooth surfaces.
Given the expansive literature on neural implicit representations, it is unrealistic to evaluate them all. Instead, we select the methods representative of new optimization paradigms: UniSURF [75] as an occupancy-grid method; Neus [109] and VolSDF [113] for jointly using volumetric and surface representations; MonoSDF [116] for integrating 3D priors in the SDF learning; iNGP [71] as a representative of fast feature-grids-based methods; NeuralAngelo [59] for integrating feature-grids into the SDF learning; BakedSDF [114] that jointly optimizes for the geometry and the texture of the scene; and the Nerfacto [100] implementation Nerf [69] that integrates several contributions, including [71], that the authors have found to work well in real-case scenarios.
@@
Benchmarking 3D reconstruction algorithms.  Existing datasets and benchmarks [98, 90, 53, 92, 86, 85, 43, 48, 9] are tailored to measure the accuracy of the 3D scene geometry itself. In contrast, we are interested in determining to what degree 3D reconstruction algorithms can be used to aid object pose estimation. Thus, we evaluate the performance of a 3D reconstruction algorithm by the accuracy of the object pose estimates it facilitates. We also evaluate how reconstruction performance (as measured by existing benchmarks) correlates to the object pose performance.
@@
Object pose benchmarks.  Pose evaluation datasets provide object CAD models, images depicting those objects in various positions and the object poses [41, 13, 44, 28, 51, 104, 111, 25, 83, 101]. These many datasets and benchmarks are unified in the common BOP benchmark by Hodan et al  [45, 47] that unifies data format, metrics, and evaluation tools. We follow the guidelines defined in the BOP benchmark to evaluate the object pose estimation. Given the ample set of datasets, we focus the benchmark on the YCB-V dataset [111], made out of the YCB dataset [17] and that encompasses objects most representative of daily-life scenarios and robotic manipulation challenges. Contrary to the BOP benchmark that focuses on the pose estimators in themselves, this benchmark focuses on the 3D reconstruction methods and how to improve them as to make the deployment of existing pose estimation algorithms easier in practice.
3 Benchmark
@@
The proposed benchmark answers the question of how well current 3D reconstruction algorithms can replace the CAD models commonly used in object pose estimation. To this end, we compare the performance of SotA object pose estimators [55, 12] when using the 3D reconstructed models against the use of classical CAD models. We follow the standard object pose evaluation guidelines defined in the BOP benchmark [45, 47] that provides accurate CAD models used as a reference for the reconstruction evaluation and to estimate the baseline object poses. The BOP benchmark also releases images of the YCB-V [111, 17] objects that could be used for 3D reconstruction but these are captured with carefully controlled light and calibrated support. Instead, we aim to evaluate the 3D reconstructions from images captured in setups close to real-life applications. This motivates the automatic capture of new images with a camera mounted on a robotic arm in an uncontrolled indoor environment. The reconstruction algorithms run on these calibrated images and the resulting meshes are compared to the reference CAD models on their use for object pose estimation. Our benchmark thus consists of two components: (1) a set of novel images of a variety of different objects with known intrinsic and extrinsic parameters that can be used to build 3D models; (2) The set of BOP test images to evaluate the meshes on their integration into the object pose estimation pipeline. The test images show the objects in different scenes and provide ground truth poses for the objects used to evaluate object pose estimation. We adopt the YCB-V [111] objects for their variety and their recurrence in the everyday life [17]. Released Data: The dataset contains calibrated and undistorted images for all 21 objects, object masks, and object reconstructions. The rest of this section describes the benchmark setup.
@@
Data Acquisition.  The images are collected with a Basler Ace camera acA2440-20gc [1] mounted on the flange of a 7 Degrees-of-Freedom (DoF) KUKA LBR IIWA 14 R820 (Supp-A). The image resolution is 2448x2048 pixels and the field of view covers the whole object. The acquisition is done in an open-floor indoor environment exposed to neon lighting and varying sunlight through ceiling-high windows, which is typical of daily life and industrial scenarios. The camera exposure remains fixed during the object’s scan. The camera and the hand-eye calibrations are run with OpenCV [15] and MoveIt [20, 4] with an error of up to 5mm in translation and 0.1 in rotation. The resulting camera calibration is used to undistort the images using COLMAP [87].
As illustrated in (Supp-Fig.1), the objects are positioned on a tripod such that the centroid of the object is approximately at the center of the reachability domain of the robot. For each object, dense outside-in views are captured by positioning the camera on a sphere centered on the object’s centroid and with a radius $\sim$30cm. The positions are evenly distributed on the sphere with latitude and longitude spanning over [0,150] and [0,360] respectively, with intervals of 10. The lowest camera position, at latitude 150 allows to capture the objects from all sides. Since the object is a source of collisions, the number of reachable views vary depending on the object’s shape and size. The number of images per object varies between 397 and 505 images with an average of 480. The motion planning and the collision avoidance are operated with the IIWA stack [38] and MoveIt [20] with the CHOMP planning algorithm [80].
4 3D Object Reconstruction
We evaluate 3D reconstructions based on the performance of the object pose estimation when replacing the traditional CAD models with the reconstructions. This section describes the benchmark setup including the evaluated reconstructions, their registration with the YCB-V [111] CAD models, the reconstruction metrics and the pose metrics, and the implementation details.
Methods.  We evaluate the MVS methods Colmap [87, 88] and Reality Capture [82]: these methods exploit feature matching and multi-view geometry to derive a sparse 3D structure of the object later used to derive depth maps [11]. The neural volumetric implicit representations Nerfacto [100] and iNGP [71] also output depth maps: a tiny MLP takes an input a position, a direction, and a learnable feature encoding to output a volume density and a color. The depth map for a given calibrated image is generated by integrating the volumetric density along rays from the camera center to the pixels [49]. Nerfacto [100] also includes adaptive sampling along the ray that typically helps the training focus on the useful parts of the scene, i.e., where the object is located. For all these methods, a mesh is obtained from the depth maps using the Poisson surface reconstruction algorithm [52].
Neural surface implicit representations train a network to approximate an SDF. UniSURF [75] approximates the SDF by an occupancy network which root values define the surface whereas Neus [109] approximates the SDF directly with a neural network to reach better reconstruction accuracy. MonoSDF [116] integrates geometric and depth cues in the SDF learning to further improve the reconstruction quality. In parallel, VolSDF [113] bridges the gap between volumetric and surface representations by training a volume density network as a function of an SDF to better constrain the learning of the scene’s geometry. BakedSDF [114] builds on top of VolSDF and optimizes a baked version of it, i.e., pre-computed values, which improves both the geometry and the texture and accelerates the rendering. NeuralAngelo [59] also addresses the runtime challenges of SDF rendering by training an SDF on learnable features represented by the hash grid introduced by the fast iNGP [71] for volumetric representations. A mesh is obtained from the SDF using the Marching Cube algorithm [64].
We also experiment with Plenoxels [30] that adopts a different representation where the scene is a feature grid with density values and spherical harmonics optimized with differentiable rendering. The absence of a neural network makes the training faster than the traditional NeRF [69] methods.
@@
The meshes thus obtained are used in place of the original YCB-V [111] CAD models to generate object renderings used in the render-and-compare pose estimators. Another way to generate rendering from implicit representations is to synthesize novel views directly from the trained network. While this has the advantage of bypassing the mesh generation [64, 52], it remains, for now, slower than physical rendering for which efficient implementations are readily available. Still, we evaluate how well can implicit renderings serve pose estimation with iNGP [71] for its fast rendering compared to the prohibitive runtime of other methods (e.g., the combination of Megapose and Nerfacto [100] takes close to 2 minutes on A100 GPU for a single image).
@@
Pose Refinement.  All reconstructions are run on the undistorted images with the camera intrinsics and extrinsics collected during the data acquisition (Sec. 3). The camera extrinsics are obtained from the kinematic chain of the robot which can be noisy so we refine them in two steps: first with the iNGP [71] pose refinement then with bundle adjustment [87] (Supp-A1). After this step, the camera extrinsics remain fixed. Even though some of the evaluated methods have pose refinement capabilities, we chose to turn them off in the main paper to focus on the evaluation of the reconstruction.
@@
Reconstruction and Texturing.  We use the default parameters for COLMAP [87, 88], which are already optimized for reconstruction purposes. Before running the Poisson surface reconstruction [52], we filter out the room’s background by roughly cropping the dense point cloud around the object. A similar processing is applied to Reality Capture [82]. The learning-based methods undergo the pre-processings described by their respective authors (see Supp-B1). The texture of learning-based methods is generated by querying the network for the color at the position of the mesh’s faces and Reality Capture [82] uses proprietary code. For methods that do not have texturing capabilities (COLMAP [87] and iNGP [71]), we evaluate the colored mesh since this leads to better pose estimates than if the mesh was textured with an off-the-shelf texturing software [106] (see Supp-C4).

### v2

@@ Abstract
Current generalizable object pose estimators, i.e., approaches that do not need to be trained per object, rely on accurate 3D models. Predominantly, CAD models are used, which can be hard to obtain in practice. At the same time, it is often possible to acquire images of an object. Naturally, this leads to the question of whether 3D models reconstructed from images are sufficient to facilitate accurate object pose estimation. We aim to answer this question by proposing a novel benchmark for measuring the impact of 3D reconstruction quality on pose estimation accuracy. Our benchmark provides calibrated images suitable for reconstruction and registered with the test images of the YCB-V dataset for pose evaluation under the BOP benchmark format. Detailed experiments with multiple state-of-the-art 3D reconstruction and object pose estimation approaches show that the geometry produced by modern reconstruction methods is often sufficient for accurate pose estimation. Our experiments lead to interesting observations: (1) Standard metrics for measuring 3D reconstruction quality are not necessarily indicative of pose estimation accuracy, which shows the need for dedicated benchmarks such as ours. (2) Classical, non-learning-based approaches can perform on par with modern learning-based reconstruction techniques and can even offer a better reconstruction time-pose accuracy tradeoff. (3) There is still a sizable gap between performance with reconstructed and with CAD models. To foster research on closing this gap, the benchmark is made available at https://github.com/VarunBurde/reconstruction_pose_benchmark.
@@ 1 Introduction
Object pose estimators predict the position and orientation of an object in a given image and is a prerequisite to various applications including robotic manipulation and Augmented Reality (AR). Existing methods can be categorized into CAD-based and CAD-free methods based on whether the pose estimation involves a CAD model, i.e., a 3D model, of the object. CAD-based methods usually lead the BOP benchmark for object pose estimation [45, 46, 2] but their deployment is limited by their reliance on CAD models. Often, such 3D models are generated by CAD applications in complex setups or reconstructed via range scanners [25] that produce highly accurate meshes. Naturally, this process can be time-consuming. Also, in many practical settings, e.g., service robots, it is infeasible to obtain such 3D models for each object class and object instance that might be encountered. A more practical approach in such scenarios would be to reconstruct 3D object models on the fly from images.
@@
Generalization is another factor limiting the deployment of pose estimators: in the example where a robot faces a new object and generates a 3D model on the fly, most of the estimators still require training or finetuning on the new object. This induces a computational overhead not only because of the network training but also because the robot must capture enough views for the training to converge. Instead, deploying CAD-based estimators that generalize to unseen objects [55, 110, 76, 16] is more efficient since they can be used without finetuning and can run with 3D reconstructions that work even with sparse views.
In this paper, we investigate using 3D reconstructions from RGB images for object pose estimation. In particular, we want to quantify the object pose accuracy gap between using high-quality CAD models and 3D models obtained from images inside generalizable object pose estimators [55, 110]. For this purpose, we propose a new benchmark. Compared to existing benchmarks for 3D reconstruction [89, 52, 47, 102], our benchmark does not treat 3D reconstruction as a task unto itself but rather evaluates the resulting 3D models inside a higher-level task (namely object pose estimation). Thus, we mainly measure 3D reconstruction performance by the accuracy of the estimated poses rather than the accuracy of the resulting 3D model itself. Results show there is no strong correlation between the two performances, i.e., one method can generate more accurate 3D models than another yet lead to less accurate pose estimates, demonstrating the need for the proposed benchmark.
@@
Our benchmark is based on the widely adopted YCB-V [112] dataset for object pose evaluation, which consists of 21 objects selected from the YCB dataset [15], including small objects, objects with low texture, complex shapes, high reflectance, and multiple symmetries. For each object, our benchmark consists of two components: (1) A new set of object images captured to be suitable for 3D reconstruction, i.e., high resolution, object-focused, and occlusion-free images contrary to the YCB-V images. (2) A set of test images for object pose estimation together with ground truth poses registered with the image sets from (1). Compared to the data for the first component, which we captured ourselves, we use the test images provided by the YCB-V [112] dataset for the second component. We believe that registering the proposed novel reconstruction benchmark with an existing pose estimation dataset that is well-known to the community will make the adoption of our benchmark easier compared to also building an object pose estimation benchmark. Given the two components, we evaluate multiple state-of-the-art 3D reconstruction methods, including both classical methods based on multi-view stereo [88, 82] and recent learning-based approaches using neural implicit representations [113, 117, 114, 75, 109]. Our experiments show that standard metrics for 3D reconstruction quality (both geometry-based and appearance-based) do not necessarily directly correlate with object pose estimation accuracy. This clearly shows a need for novel benchmarks such as ours that directly measure the impact of 3D reconstruction algorithms on object pose estimation’s accuracy.
@@
In summary, this paper makes the following contributions: (1) We propose a novel benchmark to evaluate 3D reconstruction algorithms in the context of object pose estimation. Our benchmark aims to measure to what degree 3D models reconstructed from images instead of CAD models can be used for object pose estimation. (2) We evaluate a wide range of state-of-the-art 3D reconstruction algorithms on our benchmark in terms of object pose accuracy, data requirements to reach high accuracy and reconstruction times. We draw several insights that are very relevant to the reconstruction and pose estimation research communities, including: (a) There is a clear need for benchmarks that evaluate 3D reconstruction algorithms in the context of object pose estimation. (b) Classical, non-learning-based approaches can perform similar to modern learning-based reconstruction techniques while offering faster reconstruction times. (c) The top-performing reconstruction methods are the same for the two object pose estimators considered in our benchmark [55, 110], suggesting that all object pose estimators might potentially benefit from progress in the area of 3D reconstruction. (d) While it is possible to obtain a similar performance as with CAD models for some types of objects, there is still a sizable gap in performance for other types of objects, especially for objects with fine details and reflective surfaces. Clearly, the problem of 3D object reconstruction is not yet solved. (3) The benchmark data is available online and includes calibrated images suitable for 3D reconstruction and the 3D reconstructed models.
@@ 2 Related Work
Object Pose Estimation  is addressed with multiple solutions including feature matching [27, 41, 21, 67, 84, 66, 99, 76], template matching [38, 111, 40, 39], pose regression [112, 61, 36], 3D-coordinate-regression [43, 107, 34, 11, 12, 59, 79, 103, 14], and render-and-compare [57, 54, 55, 93, 78]. One key property for the deployment of pose estimators is their generalization that falls into three categories: no generalization, generalization to object categories seen at training time [60, 108, 18, 66], and generalization to any object [55, 110, 6, 73, 76, 70, 16]. Pose estimators that generalize to any objects are the most suitable in practical scenarios as it is unrealistic to train the estimators on any object instance or category that might be encountered during deployment. However, they often rely on CAD models and it is also impractical to generate CAD models for all objects beforehand. One solution is to develop novel pose estimators that use alternative 3D representations such as Structure-from-Motion (SfM) point clouds [99, 35], implicit rendering [17, 78], and 3D models generated with implicit functions [110] or diffusion [73, 63]. The proposed benchmark follows another avenue which is how well 3D reconstruction methods measure against CAD models in their use for pose estimation. To do so, we fix the pose estimation and analyze its relative performance with various 3D reconstructions. To the best of our knowledge, there is no other benchmark that evaluates 3D reconstruction methods based on their performance on the downstream object pose estimation task. The results drawn by the proposed benchmark draw novel insights relevant for the pose estimation community on the advantages and limitations of current reconstruction methods, e.g., for which types of objects 3D reconstruction can replace CAD models. At the time of the writing, the RGBD-based FoundationPose [110] sets the state-of-the-art which motivates its integration in the benchmark to evaluate the 3D reconstructions. We also integrate Megapose [55] for it adopts a similar render-and-compare approach but on RGB images. Please refer to the supplementary for a detailed description of these methods.
@@
3D Reconstruction.  Given a set of calibrated images with known camera poses, Multi-View Stereo (MVS) approaches establish dense correspondences between images, which in turn are used to reconstruct the 3D structure of the scene. Traditionally, MVS approaches use explicit representations such as dense point clouds [31, 56], depth maps [97, 88, 10, 49], surface hulls [30, 95], or voxels [92, 53, 94, 105]. If desired, meshes can be extracted from these representations, e.g., using Poisson surface reconstruction [51]. In our benchmark, we evaluate two MVS-based 3D reconstruction approaches: the open-source COLMAP [87, 88] that commonly is used as a baseline for 3D reconstruction methods and the state-of-the-art commercial system Reality Capture [82] known for its efficiency.
@@
Alternatives to explicit reconstructions are implicit representations [28, 22, 72], i.e., continuous functions defined over the 3D space that embed properties about the scene, e.g., Signed Distance Function (SDF). Defining implicit functions can be complex so it is convenient to learn them instead [68, 77, 109, 96]. With the advent of differentiable rendering, several implicit functions relevant for 3D reconstruction can be learned simply from calibrated-image-supervision including volumetric radiance field [69], occupancy network [75], and SDF [115, 74]. Volumetric methods [69, 62] represent the scene as a colored volume where a network outputs a volume density and a color for each point in the space. Rendering the volume density [48] produces depth maps suitable to generate 3D meshes using Poisson surface reconstruction [51]. Volumetric methods are optimized for novel view synthesis only whereas optimizing SDF networks involves more geometric constraints [33] to ensure that the network exhibits SDF properties such as zero-level set corresponding to the surface.
@@
Benchmarking 3D Reconstruction Algorithms.  Existing datasets and benchmarks [98, 89, 52, 91, 85, 86, 42, 47, 8] are tailored to measure the accuracy of the 3D scene geometry itself. In contrast, we are interested in determining to what degree 3D reconstruction algorithms can be used to aid object pose estimation. Thus, we evaluate the performance of a 3D reconstruction method by the accuracy of the object pose estimates it facilitates. In parallel, we also evaluate how reconstruction performance (as measured by existing benchmarks) correlates to the object pose performance, i.e., can one really on the reconstruction performance to assess whether the reconstructed 3D model is suitable for object pose estimation. The results show that one reconstruction method that is better according to existing reconstruction benchmarks can lead to worse pose estimation than another method. This clearly shows the need for our novel benchmark, which evaluates 3D reconstruction methods based on the resulting pose accuracy.
@@
Object Pose Estimation Benchmarks.  Pose evaluation datasets provide object CAD models, images depicting those objects in various positions and the object poses [40, 11, 44, 26, 50, 104, 112, 24, 83, 101]. These many datasets and benchmarks are unified in the common BOP benchmark by Hodan et al. [45, 46] that sets the data format, metrics, and evaluation tools. We follow their guidelines to evaluate object pose estimation accuracy. Contrary to the BOP benchmark that focuses on the pose estimators, our benchmark focuses on the 3D reconstruction methods and how to improve them as to make the deployment of existing pose estimation algorithms easier in practice.
We chose to base the proposed benchmark on the YCB-V dataset [112] for two reasons: (1) it is an established dataset for object pose estimation and it displays objects with various sizes and textures relevant to the evaluation. (2) The objects shown in the dataset are available [^1] https://www.ycbbenchmarks.com/, allowing us to capture new images of them suitable for 3D reconstruction. Such images should show the objects without occlusions (to enable reconstructing the full object), focus on the object (as not to lose pixels / resolution to the background), and be of high resolution (to facilitate high-quality reconstruction). The evaluation datasets in the BOP benchmark, including YCB-V [112], provide images with the opposite properties, which makes for interesting pose estimation challenges, but the images are unsuitable for high-quality 3D reconstruction. One exception is the T-LESS [44] dataset but it is limited to small objects with grey uniform and Lambertian texture, which would severely limit the insights drawn by the evaluation as shown in the results. The combination of (1) and (2) ensures that we do not have to create a new benchmark for object pose estimation and instead can use an existing one to evaluate 3D reconstructions. We believe that using a dataset from an existing pose estimation benchmark that is well-known to the community will make the adoption of the proposed benchmark easier. To the best of our knowledge, there is currently no other existing dataset satisfying both conditions.
3 Dataset
@@
The images used for the reconstruction are captured with a camera mounted on a robotic arm in an indoor environment. The reconstruction algorithms run on these calibrated images and the resulting meshes are compared to the reference CAD models on their use for object pose estimation. Our benchmark thus consists of two components: (1) a set of novel images of a variety of different objects with known intrinsic and extrinsic parameters that can be used to build 3D models; (2) The set of BOP test images to evaluate the meshes on their integration into the object pose estimation pipeline. The test images show the objects in different scenes and provide ground truth poses for the objects used to evaluate object pose estimation. Released Data: The dataset (1) contains calibrated and undistorted images for all 21 objects, object masks, and object 3D reconstructions.
@@
Data Acquisition.  The images are collected with a Basler Ace camera acA2440-20gc [1] mounted on the flange of a 7 Degrees-of-Freedom (DoF) KUKA LBR IIWA 14 R820 on a sphere around the object (see Sec. A in the Supp.). The positions are evenly distributed on the sphere with latitude and longitude spanning over [0,150] and [0,360] respectively, with intervals of 10. The number of images per object varies between 397 and 505 images with an average of 480. The image resolution is 2448x2048 pixels and the field of view covers the whole object. The camera exposure remains fixed during the object scan. The camera and the hand-eye calibrations are run with OpenCV [13] and MoveIt [19, 4] with an error of up to 5mm in translation and 0.1 in rotation. The resulting camera calibration is used to undistort the images using COLMAP [87]. The camera poses obtained from the kinematic chain of the robot are later refined with an offline optimization (see Sec.4). The motion planning and the collision avoidance are operated with the IIWA stack [37] and MoveIt [19] with the CHOMP planning algorithm [80].
4 3D Object Reconstruction Benchmark
The proposed benchmark answers the question of how well can the 3D models generated by current 3D reconstruction methods replace the CAD models commonly used in object pose estimation. To this end, we compare the performance of two State-of-the-Art (SotA) object pose estimators [55, 110] when using the 3D reconstructed models against the use of classical CAD models. This section describes the benchmark setup including the evaluated pose and reconstruction methods, and the registration of the reconstructions with the YCB-V [112] CAD models.
Evaluated Pose Estimators.  In this work, we are interested in pose estimators that generalize to objects unseen during training and with high performance to facilitate the deployment of pose estimators in real-world scenarios. At the time of the writing, the state-of-the-art is led by the CAD-based FoundationPose [110] that operates in a render-and-compare fashion on RGB-D images. We also evaluate Megapose [55] for it adopts a similar render-and-compare strategy but on RGB-only images, which can make it more practical in real-world scenarios where RGB cameras are cheaper to deploy and used more frequently.
Evaluated Reconstruction Methods.  Given the expansive literature on neural implicit representations, it is unrealistic to evaluate them all, especially given the computational complexity of the benchmark. It roughly takes 8 days on 8 Nvidia A-100 GPUs to evaluate the following methods with a single pose estimator, which is a non-negligible cost. Instead, we select gold-standard methods or methods representative of new optimization paradigms. We release the data used in the benchmark and invite external contributions to extend the benchmark to more methods.
UniSURF [75] is a surface representation that approximates the SDF with an occupancy network, contrary to the majority of methods, such as NeuS [109], that approximates the SDF directly with a neural network to reach better reconstruction accuracy. MonoSDF [117] integrates off-the-shelf geometric and depth cues in the SDF learning process to further improve the reconstruction quality. In parallel, VolSDF [113] improves the geometry learned with volumetric representations by constraining the volumetric density to be a function of the scene’s surface geometry defined by an SDF. BakedSDF [114] builds on top of VolSDF to learn an implicit representation that is then baked into a triangle mesh, i.e., it is stored in the mesh, which is more efficient to render. The mesh is augmented with spherical Gaussians at each vertex that are further trained to represent the view-dependent appearance of the mesh. NeuralAngelo [58] improves the resolution of the mesh’s surface in a memory-efficient way by adopting the multi-resolution feature hash-grids introduced by iNGP [71] on top of which tiny-MLPs predict the surface field. A mesh is obtained from the SDF using the Marching Cubes algorithm [64]. We also evaluate Plenoxels [29] that represents the scene with a feature grid with density values and spherical harmonics optimized with differentiable rendering. The absence of a neural network makes the training faster than the traditional NeRF [69] methods. As for the volumetric methods, we evaluate iNGP [71] as a representative of fast feature-grids-based methods and the Nerfacto [100] implementation of Nerf [69] that integrates several contributions, including [71], that the authors have found to work well in real-case scenarios. One relevant property of Nerfacto [100] is its adaptive sampling along the ray that typically helps the training focus on the useful parts of the scene, i.e., where the object is located. Last, we evaluate the MVS methods COLMAP [88, 87] and Reality Capture [82].
@@
The meshes thus obtained are used in place of the original YCB-V [112] CAD models to generate object renderings used in the render-and-compare pose estimators. Another way to generate rendering from implicit representations is to synthesize novel views directly from the trained network. While this has the advantage of bypassing the mesh generation [64, 51], it remains, for now, slower than mesh rendering for which efficient implementations are readily available. Still, we evaluate how well can implicit renderings serve pose estimation with iNGP [71] for its fast rendering compared to the prohibitive runtime of other methods (e.g., Megapose with Nerfacto [100] takes close to 2 minutes on A100 GPU for a single image).
@@
Pose Refinement.  All reconstructions run on the undistorted images with the camera intrinsics and extrinsics collected during the data acquisition (see Sec. 3). The camera extrinsics are obtained from the kinematic chain of the robot which can be noisy, thus we refine them in two steps: first with the iNGP [71] pose refinement then with bundle adjustment [87] (see Sec. A.1). After this step, the camera extrinsics remain fixed. Although some evaluated methods have pose refinement capabilities, we turn them off so that all reconstruction methods share the same coordinate frame to avoid introducing biases in the evaluation.
@@
Reconstruction and Texturing.  We use the default parameters for COLMAP [87, 88], which are already optimized for reconstruction purposes. Before running the Poisson surface reconstruction [51], we filter out the room’s background by roughly cropping the dense point cloud around the object. A similar processing is applied for Reality Capture [82]. The learning-based methods undergo the pre-processings described by their respective authors (see Sec. B.1 in the supp. mat.). The texture of learning-based methods is generated by querying the network for the color at the position of the mesh’s faces and Reality Capture [82] uses proprietary code. For methods that do not have texturing capabilities (COLMAP [87] and iNGP [71]), we evaluate the colored mesh since this leads to better pose estimates than if the mesh was textured with an off-the-shelf texturing software [106] (see Sec. C.4 in the supp. mat.).